A water tank has an inlet pipe and a drain
pipe. A full tank can be emptied in 30 minutes if the drain
is opened and an empty tank can be filled in 45 minutes
with the inlet pipe opened. If both pipes are accidentally
opened when the tank is full, then how long will it take to
empty the tank?

Answer by stanbon(75887)

A water tank has an inlet pipe and a drain pipe.
-----------------------------------
A full tank can be emptied in 30 minutes if the drain
is opened and an empty tank can be filled in 45 minutes
with the inlet pipe opened.
--------------------------
Emptying Pipe DATA:
time = 1/2 hr/job ; rate = 2 jobs/hr
-----------------------------------
Fill tank DATA:
time = 3/4 hr/job ; rate = 4/3 job/hr
------------------
If both pipes are accidentally
opened when the tank is full, then how long
will it take to empty the tank?
Together DATA:
time = x hrs/job ; rate = 1/x job/hr.
======================================
Equation:
Empty rate - Fill rate = together rate
2 - (4/3) = 1/x
---
Multiply thru by 3x to get:
6x-4x = 3
2x = 3
x = 3/2 hrs
===============
Cheers,
Stan H.